OpenStudy (anonymous):

Your grades on four math tests are 97, 88, 79 and 92. What grade do you eed on the fifth test to reach an average of 90? What method would I use to solve this?

OpenStudy (anonymous):

97+88+79+92+X/5=90 450=97+88+79+92+X 450=356+X 94=X

OpenStudy (anonymous):

wow thank you!!